A chemist has one solution of citric acid that is 20% acid and another solution of 80% citric acid. She plans to mix these solutions together to make 200 liters of a new solution that is 50% acid. How many liters of each solution does she need to mix.
Answer by stanbon!(97)

Follow the acid.
Start with 20%x acid
Add 80%(200-x) acid
End up with 50%(200)acid
EQUATION:
20x +80(200-x) = 50(200)
20x +16000 - 80x = 10000
-60x = -6000
x = 100 liters of 20% solution
200-x= 100 liters of 80% solution
Cheers,
Stan H.